OpenStudy (anonymous):

3 sqrt(2x) + x sqrt(8x) - 5 sqrt(18x) and got -10x sqrt 2...what went wrong?

OpenStudy (anonymous):

8x=4*2x=2sqrt(2x) 18x=9*2x=3sqrt(2x)

OpenStudy (anonymous):

find x? equation=0?

OpenStudy (precal):